``Before, solving the problem, I just want to give you the heads up. It's
Systems of Equation: Application. I can't find the specific title to relate
with my story problems. My chapter in the book is Chapter Eight, Solving Systems by Substiution, Additional. Hope it helps!
*Angles A and B are complementary. Angle B is 18 greater than angle A. Find the measures of the two angles.*

Answer by Mathtut(3670)

since A and B are complementary
:
A+B=90.......eq 1
B=A+18.......eq 2
:
take B's value of A+18 from eq 2 and plug it into eq 1
:
A+A+18=90
:
2A=72
:
A=36 degrees
:
B=A+18=36+18=54 degrees